After robbing a bank in Dodge City, a robber gallops off at 14 mi/h. 10 minutes later, the marshall leaves to pursue the robber at 15 mi/h. How long (in hours) does it take the marshall to catch up to the robber?
Answer by mananth(16949)

In such problems you have to find the catchup distance and the catch up speed
Robber speed = 14 mph
he has a lead time of 10 minutes. = 1/6 hour.
when Marshall starts the robber is 1/6 * 14 = 14/6-> 7/3 miles away
Marshall speed = 15 mph
catchup speed = 15-14 = 1mph
time = d/r
t= (7/3) / 1
so Marshall will take 7/3 hours to catch him
= 2 hours and 20 minutes